Natasha can pick eight tree section of orange trees in four hours. It takes her brother five hours to pick the same section.
how long does it take them working together?
Suppose Natasha worked for two hours and then they worked together. What is the total amount of time that natasha would need to work to finish the job?

Answer by rwm(914)

so in one hour she can do 8/4=2 sections
and does 2/8= 1/4 in one hr
and brother can do 8/5=1 3/5=1.6 sections in one hr
1.6/8=1/5 of the job in one hr.
x/4+x/5=1
5x+4x=20
9x=20
x=20/9
x=2 2/9 hrs
if nat works two hrs alone then she has done half the job
which leaves 4 sections to do or 1/2 the job to do
x/4+x/5=1/2
9x=10
x=1 1/9 hrs.
so nat worked 2+1 1/9=3 1/9